Barefoot sells about 1,200 t-shirts per month at $27 each. The own price elasticity for these...

  1. Barefoot sells about 1,200 t-shirts per month at $27 each. The own price elasticity for these shirts is estimated to be 1.25, if the price of the shirts are increased to $30:
    1. How many shirts will they sell at the new price?
    2. The revenue will change by?
    3. Consumers will better or worse of as a result of the price change?
Answer #1

a) Price elasticity = 1.25

percentage change in price = [30-27/(30+27/2)] = [3/28.5] = 10.53%

Percentage change in quantity: 1.25= percentage change in Q/10.53

Percentage change in Q = 10.53*1.25 = 13.16%

The quantity will decrease by 13.16% to 1042

b) TR before = 1200*27 = 32400

TR after = 1042*30 = 31260

Revenue changes by 32400-31260 = 1140

c) Consmers are worse off as the equilibrium quantity decreases and price increases.